OpenStudy (anonymous):

Formula : T = d/500 + 1/2 Question : Solve the formula for d

OpenStudy (anonymous):

T = d/500 + 1/2 T - 1/2 = d/500 500(T - 1/2) = d 500T - 250 = d d = 500T - 250 -----------this is final answer....

OpenStudy (anonymous):

THANK YOU <3 !

OpenStudy (anonymous):

Lifesaver.

OpenStudy (anonymous):

You are welcome :)